OpenStudy (amazingbread12):

Find the rate that would be better to solve the problem. A sunflower grows 28 cm in 8 days. How many centimeters will it grow in 12 days? A.cm/day B.day/cm

OpenStudy (aliqanber):

(28/8) per day, so (28/8) * 12 = 42 for 12 days.

OpenStudy (anonymous):

so, cm/day? Or day/cm.

OpenStudy (aliqanber):

28/8 = 3.5 cm/day 8/28 = 0.286 day/cm

OpenStudy (anonymous):

so, it's day/cm?